# Angle-dependent planar thermal Hall effect by quasi-ballistic phonons in black phosphorus

### v1

@@ Abstract
The origin of the phonon thermal Hall effect in insulators is a matter of ongoing debate. The large amplitude of the signal in an elemental non-magnetic solid, such as black phosphorus (BP) calls for a minimal mechanism with no role for spin degree of freedom. Here, we show that a longitudinal heat flow generates a transverse temperature gradient in BP even when the magnetic field, the heat current and the thermal gradient lie in the same plane. The long phonon mean-free-path leaves little room for scattering by point-like symmetry breaking defects. we show that the signal peaks when the magnetic field is oriented along one of the two diagonal orientations of the puckered honeycomb plane and argue that this can be understood as the sum of two distinct contributions each parallel to a mirror plane. This angular dependence as well as the order of magnitude of the observed signal points to the torque exerted by magnetic field on electric dipoles traveling with heat-carrying phonons as the driver of the effect.
@@
The thermal conductivity is a second rank tensor linking the heat current density and the temperature gradient vectors. The thermal Hall effect (THE) refers to non-zero off-diagonal components of this tensor, $\overline{\kappa}$. Its phonon-driven origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26] interest. More recently, attention has focused on the observation of a planar thermal Hall effect [16, 17, 15]. This is a configuration with the three relevant vectors (Heat current, temperature gradient and magnetic field) lying in the same plane. The persistence of the signal when the signal is forbidden by crystal symmetry was attributed to defects breaking the local symmetry [17].
@@
Figure 1: Crystal structure, phonon mean free path and two versions of the thermal Hall effect.(a) Top (two layers) and side (single layer) view of the lattice of black phosphorus (BP). P atoms, marked in blue and red, belong to atomic sites with distinct environments. The $x$ and $z$ axes respectively correspond to the armchair and zigzag directions of the BP layer plane. (b) Longitudinal thermal conductivity (from previous work [12]) along the armchair ($\kappa_{xx}$) and the zigzag ($\kappa_{zz}$) orientation, at least four times of anisotropy can be seen. (c) The temperature dependence of mean free path of phonon ($l_{ph}$) in BP for two in-plane orientations ($J_{q}||z$ axis and $J_{q}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat. It shows an increasing behavior before 10 K with cooling and is comparable to the sample thickness around 20 K, the peak temperature of $\kappa_{ii}$. (d) The schematic for measuring the thermal Hall effect, with the magnetic field ($B$) perpendicular to the longitudinal heat current ($J_{q}$) and the transverse thermal gradient ($(\nabla T)_{\perp}$). The thermal Hall conductivity tensor is denoted as $\kappa_{ij}^{k}$. (e) The schematic for measuring planar thermal Hall effect, with the magnetic field ($B$) rotating from the orthogonal to the parallel direction of the $J_{q}$ or the $(\nabla T)_{\perp}$. Two types of planar thermal Hall conductivity tensors are denoted as $\kappa_{ij}^{i}$ and $\kappa_{ij}^{j}$ respectively.
@@
Figure 2: Planar thermal Hall data. (a) The field dependence of thermal Hall angle $\nabla T_{j}/\nabla T_{i}$ , with $J_{q}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures. This allows to quantify $\kappa_{xz}^{x}$. (b) Same for $J_{q}$ along the $x$ axis, $B$ and $(\nabla T)_{\perp}$ along the $z$ axis. This allows to quantify $\kappa_{xz}^{z}$. (c) Same for $J_{q}$ and $B$ along the $z$ axis and $(\nabla T)_{\perp}$ along the $x$ axis. This allows to quantify $\kappa_{zx}^{z}$. (d) The temperature dependence of the thermal Hall angle for three different configurations at 12 T. All measurements are performed on the same sample. Note that the thermal Hall angle, $\nabla T_{j}/\nabla T_{i}$ is almost unchanged when the orientations of $B$ rotates from $J_{q}$ to $(\nabla T)_{\perp}$. On the other hand, it changes by a factor of six times when the orientations of $J_{q}$ and $(\nabla T)_{\perp}$ permute with each other. This difference combined with the anisotropy in longitudinal thermal conductivity warrants Onsager reciprocity: $\kappa_{xz}^{z}=-\kappa_{zx}^{z}$, as seen in (e). (f) Comparison of thermal Hall conductivity $\kappa_{xz}$ with the magnetic field $B$ rotating from the orthogonal ($y$ axis, from previous work [12] with the same sample) to the parallel direction of the $J_{q}$ ($x$ axis) or the $(\nabla T)_{\perp}$ ($z$ axis).
@@
Figure 4: From atomic displacements to dipolar moments (a) Lattice distortion due to the thermal expansion. (b) Phosphorus atoms surrounding an inversion center. Their displacements associated with an acoustic mode (top) will affect the local electric dipole moments (middle), generating a net dipole moment (bottom). (c) Phonon modes are associated with waves of dipolar moments. A static magnetic field will exert a torque on the dipole moment traveling with the sound wave. (d) This torque ($\tau=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$) will survive only when the magnetic field is oriented perpendicular to the wave velocity. Orange arrows show the orientation of the Lorenz forces on the two poles of the dipole.
Here, we report on the observation of a planar Hall effect in black phosphorus(P), an elemental insulator with ballistic phonons [27] and no magnetism. The thermal conductivity tensor has unequal diagonal components ($\kappa_{zz}\simeq 5\kappa_{xx}$). Nevertheless, the off-diagonal components match each other ($\kappa_{xz}(B)\simeq\kappa_{zx}(-B)$), as expected by Onsager reciprocity. We quantify the variation of the thermal Hall signal as a function of the in-plane orientation of the magnetic field and find that it displays a twofold oscillation with minimum and maximum along one of the two diagonals of the $xz$ plane of the puckered honeycomb lattice. This indicates that the signal is the sum of two sinusoidal contributions along two high-symmetry axes. A quantitative account of our observation is missing. Nevertheless, we argue that a scenario in which electric dipole moments, generated by the thermal gradient and the atomic vibrations, play a key role is compatible with the order of magnitude of the observed signal as well as its angular variation.
@@
Fig. 2a-c shows the planar thermal Hall data for three different configurations. In each panel, the orientations of the three relevant vectors (B, $J_{q}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla T_{j}/\nabla T_{i}$) is shown.
@@
The thermal conductivity tensor $\overline{\kappa}$, is thus the thermal diffusivity tensor multiplied by a scalar: $\overline{\kappa}=\overline{D}C$. Therefore, off-diagonal components in $\overline{\kappa}$ are proportional to their counterparts in $\overline{D}$. The latter is the product of carrier velocity and carrier mean free path. Sound velocity, $v_{s}$, and phonon mean free path $\ell_{ph}$ (see Fig. 1c) are both anisotropic at zero magnetic field. The issue is to find a way for magnetic field to skew one or both of these tensors off the symmetry axes
@@
Absent magnetism, the interplay between electric polarization and phonons deserves scrutiny. The computed map of charge distribution is strongly orientational [36]. Dipole-active phonon modes has been detected by infrared spectroscopy [37]. These are optical phonons and not heat-carrying acoustic modes which interest us. Nevertheless, any phonon breaking the inversion symmetry will generate an electric dipolar wave. This is a feature highlight in the context of ferrons, the elementary excitation of a ferroelectric solid, even in a paraelectric insulators [38, 39].
The thermal expansion of black P is not only anisotropic, but has opposite signs along armchair and zigzag crystalline orientations [40]. As a consequence, in presence of a temperature gradient each tetragonal unit cell is slightly distorted. Its lattice parameter, compared to its colder or warmer neighboring cells, is longer and shorter along opposite (Fig. 4a) orientations. This paves the way for finite dipolar polarization, and for chirality in presence of the magnetic field. The inversion center of the pristine unit cell is not an atomic site and atomic displacements associated with phonons generate electric dipole moments (Fig. 4b). Therefore, acoustic phonons breaking inversion centers are traveling electric dipolar moments capable of coupling to a static magnetic field (Fig. 4c).
@@
Experimentally, the thermal Hall angle at $B$ = 10 T peaks to ×$\approx\times 10^{-2}$ in black P. The length scale extracted from this angle and fundamental constants ($\lambda_{tha}$ = $\ell_{B}\sqrt{\kappa_{ij}/\kappa_{jj}}$) is about 5 Å. Intriguingly, in all insulators displaying a thermal Hall signal, this length scale lies in a narrow range (2 Å${}<\lambda_{tha}<7$ [12, 14]. As seen in figure S2 in the supplement [30], this remains true for recently studied cases of Si, Ge and quartz [15].
@@
This expectation is of the order of magnitude of the measured signal. A rigorous treatment would presumably include in this expression the Grüneisen parameter [42]. Possibly, the atomic electric dipole polarizability [43] would also be present. However, the former, which is dimensionless remains of the order of unity and the latter introduces a length scale which does not vary widely among different solids. Therefore, a scenario attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves emerges promising from our study.
@@
Supplementary Materials for “Angle-dependent planar thermal Hall effect in black phosphorus”
@@
The longitudinal ($\nabla T_{i}$) and the transverse ($\nabla T_{j}$ ) thermal gradient generated by a longitudinal thermal current $J_{q}$ were measured. They lead to the longitudinal ($\kappa_{ii}$) and the transverse ($\kappa_{ij}$) thermal conductivity, as well as the thermal Hall angle ($\nabla T_{j}/\nabla T_{i}$):
$\kappa_{ii}=\frac{Q_{i}}{\nabla T_{i}}$ (S1)
$\frac{\nabla T_{j}}{\nabla T_{i}}=\frac{\kappa_{ij}}{\kappa_{jj}}$ (S2)
$\kappa_{ij}=\frac{\nabla T_{j}}{\nabla T_{i}}\cdot\kappa_{jj}$ (S3)
@@
Fig. S1 shows the field dependent thermal Hall angles ($\nabla T_{j}/\nabla T_{i}$) with the $J_{q}$ alongs $x$ axis and the $(\nabla T)_{\perp}$ along $z$ axis at six different field orientations. $\phi$ is the angle of the magnetic field with respect to the $J_{q}$ ($x$ axis). $\phi=0^{\circ}$ and $\phi=90^{\circ}$ represent the orientations along two high-symmetry axes, and their $\nabla T_{j}/\nabla T_{i}$ signal have the same amplitude but opposite signs. $\phi=-45^{\circ}$ and $\phi=45^{\circ}$ represent two diagonal orientations of the puckered honeycomb plane that have the maximum and vanished responses respectively, suggesting that the planar thermal Hall signal consists of two distinct contributions each along one high-symmetry axis. $\phi=-45^{\circ}$ and $\phi=135^{\circ}$ have signals with opposite signs, as expected from the antisymmetric operation ($\kappa_{ij}$(B) = -$\kappa_{ij}$(-B)).
Figure S1: Raw data of angular dependent planar thermal Hall effect in Black phosphorus. Field dependent thermal Hall angle ($\nabla T_{z}/\nabla T_{x}$) at a constant temperature and six different field orientations.
Figure S2: Thermal Hall angle in different insulators The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source:  [1, 2, 3, 4, 5, 6, 8, 9, 11, 12, 14, 15]).
@@
In our experiment, the magnetic field was rotated in the $xz$ plane, the longitudinal thermal gradient thermal gradient was applied along the $x$-axis. Putting $B_{y}=0$ and $P_{y}=0$, Equation S6 becomes:
@@
One can see that the expected torque along the $z$-axis has two components. Each of them vanishes when the field is along either the $x$-axis or the $z$-axis. This is in agreement with the experimental data seen in Fig. 3 of the main text.

### v2

@@ Abstract
The origin of the phonon thermal Hall effect in insulators is a matter of ongoing debate. The large amplitude of the signal in an elemental non-magnetic solid, such as black phosphorus (BP) calls for a minimal mechanism with no role for spin degree of freedom. Here, we show that a longitudinal heat flow generates a transverse temperature gradient in BP even when the magnetic field, the heat current and the thermal gradient lie in the same plane. The long phonon mean-free-path leaves little room for scattering by point-like symmetry breaking defects. We argue that the angular dependence of the signal can be accounted for as a sum of two sinusoidal components each peaking when the magnetic field is parallel to a high symmetry axis. Anharmonicity and finite thermal expansion emerge as indispensable ingredients for a planar Hall signal. We identify the torque exerted by magnetic field on traveling electric dipolar waves as a possible microscopic driver.
@@
The thermal conductivity is a second rank tensor linking the heat current density and the temperature gradient vectors. The thermal Hall effect (THE) refers to non-zero off-diagonal components of this tensor, $\overline{\kappa}$. Its origin in insulators has attracted much recent experimental [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretical [18, 19, 20, 21, 22, 23, 24, 25, 26] interest.
The planar thermal Hall effect refers to a configuration in which the three relevant vectors (heat current, temperature gradient and magnetic field) lie in the same plane. This was first observed in the Kitaev spin liquid candidate $\alpha$-RuCl3 [4, 27, 28, 29]. More recently, it was also observed in other solids, where the thermal Hall signal is attributed to phonons [30, 31, 32]. The persistence of the signal, when it is forbidden by crystal symmetry, was attributed to defects breaking the local symmetry [31].
@@
Figure 1: Crystal structure, phonon mean free path and two versions of the thermal Hall effect.(a) Top (two layers) and side (single layer) view of the lattice of black phosphorus (BP). P atoms, marked in blue and red, belong to atomic sites with distinct environments. (b) Longitudinal thermal conductivity (from previous work [15]) along the armchair ($\kappa_{xx}$) and the zigzag ($\kappa_{zz}$) orientation, at least four times of anisotropy can be seen. (c) The temperature dependence of mean free path of phonon ($\ell_{ph}$) in BP for two in-plane orientations ($J_{q}||z$ axis and $J_{q}||x$ axis ), extracted from the longitudinal thermal conductivity, the sound velocity and the specific heat. It shows an increasing behavior before 10 K with cooling and is comparable to the sample thickness around 20 K, the peak temperature of $\kappa_{ii}$. (d) The schematic for measuring the thermal Hall effect, with the magnetic field ($B$) perpendicular to the longitudinal heat current ($J_{q}$) and the transverse thermal gradient ($(\nabla T)_{\perp}$). The thermal Hall conductivity tensor is denoted as $\kappa_{ij}^{k}$. (e) The schematic for measuring planar thermal Hall effect, with the magnetic field ($B$) rotating from the orthogonal to the parallel direction of the $J_{q}$ or the $(\nabla T)_{\perp}$. Two types of planar thermal Hall conductivity tensors are denoted as $\kappa_{ij}^{i}$ and $\kappa_{ij}^{j}$ respectively.
@@
Figure 2: Planar thermal Hall data. (a) The field dependence of thermal Hall angle $\nabla_{j}T/\nabla_{i}T$ , with $J_{q}$ and $B$ along the $x$ axis, $(\nabla T)_{\perp}$ along the $z$ axis, at three typical temperatures. This allows to quantify $\kappa_{xz}^{x}$. (b) Same for $J_{q}$ along the $x$ axis, $B$ and $(\nabla T)_{\perp}$ along the $z$ axis. This allows to quantify $\kappa_{xz}^{z}$. (c) Same for $J_{q}$ and $B$ along the $z$ axis and $(\nabla T)_{\perp}$ along the $x$ axis. This allows to quantify $\kappa_{zx}^{z}$. (d) The temperature dependence of the thermal Hall angle for three different configurations at 12 T. All measurements are performed on the same sample. Note that the thermal Hall angle, $\nabla_{j}T/\nabla_{i}T$ is almost unchanged when the orientations of $B$ rotates from $J_{q}$ to $(\nabla T)_{\perp}$. On the other hand, it changes by a factor of six times when the orientations of $J_{q}$ and $(\nabla T)_{\perp}$ permute with each other. This difference combined with the anisotropy in longitudinal thermal conductivity warrants Onsager reciprocity: $\kappa_{xz}^{z}=-\kappa_{zx}^{z}$, as seen in (e). (f) Comparison of thermal Hall conductivity $\kappa_{xz}$ with the magnetic field $B$ rotating from the orthogonal ($y$ axis, from previous work [15] with the same sample) to the parallel direction of the $J_{q}$ ($x$ axis) or the $(\nabla T)_{\perp}$ ($z$ axis).
@@
Figure 4: From atomic displacements to dipolar moments (a) Because of thermal expansion, the orthorhombic unit cell is distorted. A rectangle with two perpendicular mirrors is replaced by a trapezoid with no reflection symmetry. (b) Phosphorus atoms surrounding an inversion center. Their displacements associated with an acoustic mode (top) will affect the local electric dipole moments (middle), generating a net dipole moment (bottom). (c) Phonon modes are associated with waves of dipolar moments. A static magnetic field will exert a torque on the dipole moment traveling with the sound wave. (d) This torque ($\tau=\overrightarrow{P}\times(\overrightarrow{B}\times\overrightarrow{v})$) will survive only when the magnetic field is oriented perpendicular to the wave velocity. Green arrows show the orientation of the Lorenz forces on the two poles of the dipole.
Here, we report on the observation of a planar Hall effect in black phosphorus(P), an elemental insulator with ballistic phonons [33]. The thermal conductivity tensor has unequal diagonal components ($\kappa_{zz}\simeq 5\kappa_{xx}$). Nevertheless, the off-diagonal components match each other ($\kappa_{xz}(B)\simeq\kappa_{zx}(-B)$), as expected by Onsager reciprocity. We quantify the variation of the thermal Hall signal as a function of the in-plane orientation of the magnetic field and find that it displays a twofold oscillation with minimum and maximum along one of the two diagonals of the $xz$ plane of the puckered honeycomb lattice. This indicates that the signal is the sum of two sinusoidal contributions along two high-symmetry axes. A quantitative account of our observation is missing. Nevertheless, we argue that because of the finite thermal expansion, a temperature gradient allows the emergence of a signal forbidden in thermodynamic equilibrium. Electric dipole moments travelling with sound can couple to magnetic field and generate a thermal Hall signal in planar configuration.
@@
Fig. 2a-c shows the planar thermal Hall data for three different configurations. In each panel, the orientations of the three relevant vectors ($B$, $J_{q}$ and $(\nabla T)_{\perp}$) are sketched and the field dependence of the ratio of the transverse to longitudinal temperature gradients ($\nabla_{j}T/\nabla_{i}T$) is shown.
@@
The thermal conductivity tensor $\overline{\kappa}$, is thus the thermal diffusivity tensor multiplied by a scalar: $\overline{\kappa}=\overline{D}C$. Therefore, off-diagonal components in $\overline{\kappa}$ are proportional to their counterparts in $\overline{D}$. The latter is the product of carrier velocity and carrier mean free path. Sound velocity, $v_{s}$, and phonon mean free path $\ell_{ph}$ (see Fig. 1c) are both anisotropic at zero magnetic field. The issue is to find a way for magnetic field to skew one or both of these tensors off the symmetry axes.
@@
Absent magnetism, the interplay between electric polarization and phonons deserves scrutiny. The computed map of charge distribution is strongly orientational [42]. Dipole-active phonon modes has been detected by infrared spectroscopy [43]. These are optical phonons and not heat-carrying acoustic modes which interest us. Nevertheless, any phonon breaking the inversion symmetry will generate an electric dipolar wave. This feature highlighted in the context of ferrons, the elementary excitation of a ferroelectric solid [44, 45], persists even in a paraelectric insulator.
Black P has an anisotropic thermal expansion with opposite signs along armchair and zigzag crystalline orientations [46]. In presence of a temperature gradient, each unit cell is slightly distorted (Fig. 4a). This feature, driven by unavoidable anharmonicity in any solid [47], is the fundamental reason that a signal forbidden in thermodynamic equilibrium is allowed in presence of a temperature gradient. It also paves the way for chirality in presence of the magnetic field. The inversion center of the pristine unit cell is not an atomic site and atomic displacements associated with phonons generate electric dipole moments (Fig. 4b). Therefore, acoustic phonons breaking inversion centers are traveling electric dipolar moments capable of coupling to a static magnetic field (Fig. 4c).
@@
Experimentally, the thermal Hall angle at $B$ = 10 T peaks to $\approx 10^{-2}$ in black P. The length scale extracted from this angle and fundamental constants ($\lambda_{tha}$ = $\ell_{B}\sqrt{\kappa_{ij}/\kappa_{jj}}$) is about 5 Å. Intriguingly, in all insulators displaying a thermal Hall signal, this length scale lies in a narrow range (2 Å${}<\lambda_{tha}<7$ [15, 17]. As seen in Fig. S2 in the supplement [36], this remains true for recently studied cases of Si, Ge and quartz [32].
@@
This is of the order of magnitude of the measured signal. A rigorous treatment would presumably include the Grüneisen parameter [47] and the atomic electric dipole polarizability. The former, which is dimensionless, remains of the order of unity and the latter introduces a length scale which does not vary widely among different solids [49]. Thus, a scenario attributing the thermal Hall signal to the interaction between a static magnetic field and traveling electric dipolar waves, which to the best of our knowledge, was not considered before, emerges as a promising candidate.
Acknowledgements.
@@
Supplementary Materials for “Angle-dependent planar thermal Hall effect by quasi-ballistic phonons in black phosphorus” by X. Li et al.
@@
The longitudinal ($\nabla_{i}T$) and the transverse ($\nabla_{j}T$) thermal gradient generated by a longitudinal thermal current $J_{q}$ were measured. They lead to the longitudinal ($\kappa_{ii}$) and the transverse ($\kappa_{ij}$) thermal conductivity, as well as the thermal Hall angle ($\nabla_{j}T/\nabla_{i}T$):
$\kappa_{ii}=\frac{Q_{i}}{\nabla_{i}T}$ (S1)
$\frac{\nabla_{j}T}{\nabla_{i}T}=\frac{\kappa_{ij}}{\kappa_{jj}}$ (S2)
$\kappa_{ij}=\frac{\nabla_{j}T}{\nabla_{i}T}\cdot\kappa_{jj}$ (S3)
@@
Fig. S1 shows the field dependent thermal Hall angles ($\nabla_{j}T/\nabla_{i}T$) with the $J_{q}$ alongs $x$ axis and the $(\nabla T)_{\perp}$ along $z$ axis at six different field orientations. $\phi$ is the angle of the magnetic field with respect to the $J_{q}$ ($x$ axis). $\phi=0^{\circ}$ and $\phi=90^{\circ}$ represent the orientations along two high-symmetry axes, and their $\nabla_{j}T/\nabla_{i}T$ signal have the same amplitude but opposite signs. $\phi=-45^{\circ}$ and $\phi=45^{\circ}$ represent two diagonal orientations of the puckered honeycomb plane that have the maximum and vanished responses respectively, suggesting that the planar thermal Hall signal consists of two distinct contributions each along one high-symmetry axis. $\phi=-45^{\circ}$ and $\phi=135^{\circ}$ have signals with opposite signs, as expected from the antisymmetric operation ($\kappa_{ij}$(B) = -$\kappa_{ij}$(-B)).
Figure S1: Raw data of angular dependent planar thermal Hall effect in black phosphorus. Field dependent thermal Hall angle ($\nabla_{z}T/\nabla_{x}T$) at a constant temperature and six different field orientations.
Figure S2: Thermal Hall angle in different insulators. The transverse thermal conductivity divided by magnetic field as a function of longitudinal thermal conductivity in different insulators (source: [1, 50, 6, 7, 8, 9, 11, 12, 14, 15, 17, 32]). Note that magnetic insulators with a non-linear field dependence of the thermal Hall effect, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.
@@
In our experiment, the magnetic field was rotated in the $xz$ plane, the longitudinal thermal gradient was applied along the $x$-axis. Putting $B_{y}=0$ and $P_{y}=0$, Equation S6 becomes:
@@
One can see that the expected torque along the $z$-axis has two components. Each of them vanishes when the field is along either the $x$-axis or the $z$-axis. This is in agreement with the experimental data seen in Figure 3 of the main text.
S5 The universality in thermal Hall physics
Figure S2 illustrates the thermal Hall angle across various insulators. It presents the ratio of transverse thermal conductivity to magnetic field as a function of longitudinal thermal conductivity for these insulators. It is important to note that the figure exclusively features insulators exhibiting a linear or sub-linear relationship between transverse thermal conductivity and magnetic field. Magnetic insulators with a non-linear dependence, such as Lu2V2O7[2] and Cu(1-3, bdc)[3], are not included in the figure.